# Laura Barros        Aiko Schmeißer

### v3

@@ 5 Effects of Unemployment Insurance on Political Entry
The main results presented in the previous section show that job loss significantly increases individuals’ likelihood of joining a party and running for local office. As discussed in Section 2.3, these effects may be explained by two mechanisms: income losses that reduce the opportunity costs of political entry, and increased time resources that allow individuals to engage more intensively in politics. Both channels may be influenced by individuals’ eligibility for unemployment insurance benefits. In Brazil, workers who are dismissed without just cause can receive UI benefits for 3 to 5 months, with an average replacement rate of about 80%. [^35] The length of UI payments depends on the length of employment in the 36 months before the layoff, and the replacement rate depends on the previous earnings (varying between 100% for individuals earning the minimum wage and 67% at the benefit cap, which is at 2.65 times the minimum wage). Benefit receipt mitigates the income losses from layoffs and may thereby reduce the financial incentives to pursue a political position. At the same time, during benefit receipt dismissed workers have lower incentives to find reemployment. [^36] The negative effects of UI eligibility on formal employment have been widely documented, including for Brazil (Gerard et al., 2020; Gerard and Gonzaga, 2021; Britto et al., 2022b). Britto (2022) also shows that an extension of the UI eligibility period in Brazil decreased total employment when taking into account the positive effect on informal employment. The negative employment effects of UI eligibility decrease labor earnings, which may offset the benefit payments, and increase individuals’ time availability. The increased time resources, in turn, may facilitate individuals’ political engagement. In this section, we investigate whether eligibility for UI benefits affects laid-off individuals’ decision to enter politics, which can help shed light on the mechanisms driving the effect of job loss on political entry.
@@ 5.1 Empirical strategy
where $Y_{i}$ is an outcome of worker $i$. $X_{i}$ denotes the time since the last layoff (re-centered around the 16-month cutoff), and $D_{i}$ is a dummy for being eligible for UI benefits, i.e., $D_{i}=1(X_{i}\geq 0)$. $\beta$ is the coefficient of interest that identifies the effect of UI eligibility. [^37] As we lack information on UI take-up, our focus is on benefit eligibility. Consequently, our estimates should be interpreted as an intent-to-treat effect. According to Britto et al. (2022b), the probability of UI take-up jumps by about 60 percentage points at the 16-month eligibility cutoff. To ensure that eligible and non-eligible workers are comparable, our main results are based on a local linear model with a narrow bandwidth of 60 days around the cutoff. [^38] This choice follows the main specifications in Britto et al. (2022b), Britto et al. (2022a), and Amorim et al. (2025). In Appendix Tables C.1 and C.2 we show that our main results are robust to alternative polynomial orders and bandwidth choices, including the optimal bandwidth proposed by Calonico et al. (2014). In Appendix Figure C.4 we also perform permutation tests that compare our main estimates to the distribution of RD estimates obtained from placebo cutoff points.
@@
As in the previous section, the estimation sample consists of all workers aged 25 to 50 displaced from a full-time, open-ended, private-sector job of at least six months tenure. To increase statistical power, we extend our layoff period to cover the years 2006 to 2014.393939We start in 2006 because this allows us to observe past layoffs up to 24 months prior to the current layoff (our employment data begin in 2004) and to measure placebo political outcomes in the two cycles prior to the current layoff (our political data begin in the 2000 cycle). We stop in 2014 because of the numerous changes introduced to the UI system after that year. This allows us to measure political outcomes over two election cycles after the layoff and labor market outcomes over up to 4 years after the layoff. [^40] While we must restrict the sample to individuals with unique names for the party membership outcome, we do not impose this restriction when analyzing candidacies, which are less common and thus require larger sample sizes. In addition, we exclude from the sample all workers who were displaced on the first or the last day of each month in their last layoff. Appendix Figure C.1 shows that the number of dismissals is higher at the very beginning and the very end of the month. Dropping these observations ensures that our RD cutoff does not coincide with spikes in the density of dismissals that occur every 30 days, i.e., that are not specific to the 16-month cutoff.
@@
Appendix Figure C.2 shows that, when using the restricted sample, there is no evidence of discontinuity in the density of the running variable at the 16-month cutoff. We also confirm the continuity using the bias-robust test by Cattaneo et al. (2018, 2020). In addition, in Appendix Figure C.3 we show that individuals around the cutoff are similar in a rich set of pre-determined covariates, including gender, education, age, earnings, tenure, and sector shares. Overall, these results provide strong support for the continuity assumption required by the RD approach.
@@ 5.2 Results
Results from model (4) are presented in Table 3, and the corresponding RD plots are shown in Figure 8. As a first step, we estimate the effects of UI eligibility on labor market outcomes (Panel A, columns (1) and (2) of Table 3). We find a significant negative effect on employment that amounts to 0.68 fewer months employed over the four years after the layoff. Moreover, labor earnings decrease by 2,192 BRL. The drop in labor earnings is likely compensated by the UI payments that eligible individuals receive. [^41] Britto et al. (2022b) find that individuals barely meeting the 16-months cutoff receive additional UI benefits of about 2,100 BRL. Thus, the results suggest that UI eligibility leads to a rise in time availability following the job loss without strongly affecting the total available income.
(1) (2) (3) (4) (5) (6) Outcome: Months Labor New party membership (%) Candidacy (%) employed earnings before layoff after layoff before layoff after layoff [A] Full sample $\beta$0.679∗∗∗ -2,192.5∗∗∗ -0.002 0.150∗∗ -0.011 0.028 (0.041) (149.9) (0.080) (0.072) (0.013) (0.019) Control mean 26.138 56,281.8 4.330 3.467 0.236 0.449 Relative effect -2.6% -3.9% -0.1% 4.3% -4.6% 6.2% Observations 2,048,674 2,048,674 1,047,683 1,047,683 2,048,674 2,048,674 [B] Low educated ($<$ 12 years of education) $\beta$0.617∗∗∗ -1,822.6∗∗∗ 0.071 0.066 -0.015 -0.006 (0.057) (166.1) (0.121) (0.109) (0.019) (0.026) Control mean 25.029 47,526.6 4.264 3.454 0.221 0.428 Relative effect -2.5% -3.8% 1.7% 1.9% -6.9% -1.4% Observations 986,755 986,755 454,665 454,665 986,755 986,755 [C] High educated ($\geq$ 12 years of education) $\beta$0.708∗∗∗ -2,311.2∗∗∗ -0.059 0.216∗∗ -0.007 0.061∗∗ (0.059) (240.6) (0.106) (0.097) (0.019) (0.027) Control mean 27.143 64,213.1 4.381 3.476 0.250 0.468 Relative effect -2.6% -3.6% -1.3% 6.2% -2.7% 13.0% Observations 1,061,919 1,061,919 593,018 593,018 1,061,919 1,061,919 Note: The table presents results from model (4) for the effect of unemployment benefit eligibility on the number of months employed and the total formal labor earnings in the four years after the layoff (columns (1) and (2)), on the likelihood of newly registering as a party member in the two election cycles before and after the layoff (columns (3) and (4)), and on the likelihood of running for local councilor in the two election cycles before and after the layoff (columns (5) and (6)). The sample includes all workers displaced within a bandwidth of 60 days around the cutoff date that determines benefit eligibility, which is 16 months since the last layoff. Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means in columns (3) to (6) have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
A. Full sample Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption B. Low educated ($<$ 12 years of education) Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption C. High educated ($\geq$ 12 years of education) Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption Note: The figure plots the average outcomes of workers around the 16-month cutoff date for eligibility for unemployment benefits, for different groups of workers. Outcomes are the number of months employed in the four years after the layoff, the probability of newly registering as a party member in the two election cycles after the layoff (in %), and the probability of running for local councilor in the two election cycles after the layoff (in %). Panel A shows results among all workers in the sample, and Panels B and C distinguish workers with less than vs. at least 12 years of education. Dots show averages for 5-day bins. The lines represent a local linear polynomial smoothing using a 60-day bandwidth, together with 95% confidence intervals.
@@
Given that earlier we have documented differential effects of job loss by education, Panels B and C of Table 3 also show the effects of UI benefits separately for low-educated workers ($<$ 12 years of schooling) and high-educated workers ($\geq$ 12 years of schooling). We find positive effects on party memberships and candidacies among the highly educated but no significant impact among the low educated. The effects on workers with high levels of education are statistically significant and meaningful in size, indicating a 6.2% increase in the likelihood of becoming a party member and a 13.0% increase in the likelihood of running for office (both relative to the control mean). These findings are consistent with the time-resources mechanism: by reducing reemployment incentives, eligibility for UI benefits frees up time that can be redirected toward political engagement. As suggested by our theoretical framework in Section 2.3, the importance of this channel increases with the expected benefits of political participation. These benefits may vary with education – for example, if more educated individuals have greater chances of electoral success or are more likely to leverage political connections for gaining access to public-sector jobs. Such mechanisms can explain why both job loss and UI eligibility have stronger effects on political entry among the highly educated.
@@ 6.1 Individual-level evidence
To explore the impact of layoffs on political selection, we examine the individual characteristics of candidates and political members who enter politics as a result of being laid off. We use a combination of sample moments in an instrumental variable framework to estimate the average characteristics of three groups: (1) compliers—individuals who enter politics because of the layoff, (2) always-takers—those who enter politics regardless of whether they are laid off or not, and (3) never-takers—those who do not enter politics regardless of their layoff status. To identify compliers, we build on the kappa weighting approach introduced by Abadie (2003). We estimate the expected value of covariate $X_{i}$ for compliers as
@@
Here, $Y_{i}$ indicates whether an individual enters politics in a post-layoff cycle (either as a party member or a councilor candidate), and $Treat_{i}$ represents their layoff status. We estimate the propensity score, $P(X_{i})=P(Treat_{i}=1|X_{i})$, using the set of all covariates considered in the complier analysis. [^43] Słoczyński et al. (2025) provide a comprehensive treatment of various kappa weighting estimators that are implemented in the Stata command kappalate. We use the estimator based on their $\kappa_{i1}$ definition. The authors recommend these kappa weights for applications without always-takers (one-sided noncompliance) to avoid near-zero denominators. Similarly, in our case, the number of always-takers is very small relative to the number of never-takers.
@@
When analyzing the characteristics of political entrants, we focus on measures that capture individual competence, in addition to socio-demographic characteristics. First, we consider years of education as a standard measure of competence. [^44] For Brazil, Melo and Tigre (2024) provide evidence that more educated mayors are more effective managers of public resources, as they commit fewer municipal administration infringements. However, they are not found to engage in fewer corrupt practices or improve the provision of local public goods. Second, we examine individuals’ log hourly wage before the layoff, which, following Dal Bó et al. (2013), can be treated as a measure of skill, as the private sector typically rewards highly skilled workers with higher wages. Third, we consider person wage effects from an AKM model that regresses log wages on person fixed effects and firm fixed effects (Abowd et al., 1999). The person fixed effects measure the wage premia that workers receive independent of the firm that they work at, likely reflecting workers’ ability that is rewarded across all firms. [^45] To our knowledge, we are the first to use AKM effects to study the competence of politicians. The AKM model is fitted on all RAIS contracts in the pre-layoff years 2004 to 2008. Appendix Table D.1 reports a summary of the AKM estimation results. Moreover, we consider residualized AKM person effects, controlling for gender, race, age deciles, education categories, as well as the full set of interactions between these variables. The resulting residuals capture wage premia relative to observably similar workers. [^46] Besley et al. (2017) and Dal Bó et al. (2017) also consider residuals from Mincer wage regressions. Our AKM measures additionally purge wages from firm variation in wage-setting, which may not reflect individual ability. To validate our competence measures, we correlate them with local councilors’ misconduct detected during the electoral process and with their electoral performance. Results are reported in Appendix Table D.2 and show that all our competence measures are significantly negatively correlated with the probability of misbehaving in the electoral process and significantly positively related the the probability of being elected. [^47] The misbehavior variable captures a broad range of irregularities detected by the Brazilian Electoral Court during the elections in 2016 and 2020 that resulted in candidacy annulment. These include abuse of power, failure to meet registration requirements, vote buying, prohibited conduct, ineligibility under the “Clean Record Act” (Lei da Ficha Limpa), illicit use of campaign resources, objections to the candidacy, rejection by the party or coalition, and invalidation of the party. Finally, in addition to the competence measures, we also study socio-demographic representation in terms of gender and age.
Figure 9 presents the results. Overall, we find that individuals who enter politics because of a layoff (compliers) tend to be positively selected in terms of competence relative to always-takers and never-takers. For example, layoff-induced councilor candidates have an average of 12.0 years of education, compared to about 10.9 years among individuals who always or never run for local

### v4

@@ 5 Effects of Unemployment Insurance on Political Entry
The main results presented in the previous section show that job loss significantly increases individuals’ likelihood of joining a party and running for local office. As discussed in Section 2.3, these effects may be explained by two mechanisms: income losses that reduce the opportunity costs of political entry, and increased time resources that allow individuals to engage more intensively in politics. Both channels may be influenced by individuals’ eligibility for unemployment insurance benefits. In Brazil, workers who are dismissed without just cause can receive UI benefits for 3 to 5 months, with an average replacement rate of about 80%. [^38] The length of UI payments depends on the length of employment in the 36 months before the layoff, and the replacement rate depends on the previous earnings (varying between 100% for individuals earning the minimum wage and 67% at the benefit cap, which is at 2.65 times the minimum wage). Benefit receipt mitigates the income losses from layoffs and may thereby reduce the financial incentives to pursue a political position. At the same time, during benefit receipt dismissed workers have lower incentives to find reemployment. [^39] The negative effects of UI eligibility on formal employment have been widely documented, including for Brazil (gerard2020; gerard2021; Britto2022). britto2022_restat also shows that an extension of the UI eligibility period in Brazil decreased total employment when taking into account the positive effect on informal employment. The negative employment effects of UI eligibility decrease labor earnings, which may offset the benefit payments, and increase individuals’ time availability. The increased time resources, in turn, may facilitate individuals’ political engagement. In this section, we investigate whether eligibility for UI benefits affects laid-off individuals’ decision to enter politics, which can help shed light on the mechanisms driving the effect of job loss on political entry.
@@ 5.1 Empirical strategy
where $Y_{i}$ is an outcome of worker $i$. $X_{i}$ denotes the time since the last layoff (re-centered around the 16-month cutoff), and $D_{i}$ is a dummy for being eligible for UI benefits, i.e., $D_{i}=1(X_{i}\geq 0)$. $\beta$ is the coefficient of interest that identifies the effect of UI eligibility. [^40] As we lack information on UI take-up, our focus is on benefit eligibility. Consequently, our estimates should be interpreted as an intent-to-treat effect. According to Britto2022, the probability of UI take-up jumps by about 60 percentage points at the 16-month eligibility cutoff. To ensure that eligible and non-eligible workers are comparable, our main results are based on a local linear model with a narrow bandwidth of 60 days around the cutoff. [^41] This choice follows the main specifications in Britto2022, Britto2022a, and Amorim2023. In Appendix Tables C.1 and C.2 we show that our main results are robust to alternative polynomial orders and bandwidth choices, including the optimal bandwidth proposed by calonico2014. In Appendix Figure C.4 we also perform permutation tests that compare our main estimates to the distribution of RD estimates obtained from placebo cutoff points.
@@
As in the previous section, the estimation sample consists of all workers aged 25 to 50 displaced from a full-time, open-ended, private-sector job of at least six months tenure. To increase statistical power, we extend our layoff period to cover the years 2006 to 2014.424242We start in 2006 because this allows us to observe past layoffs up to 24 months prior to the current layoff (our employment data begin in 2004) and to measure placebo political outcomes in the two cycles prior to the current layoff (our political data begin in the 2000 cycle). We stop in 2014 because of the numerous changes introduced to the UI system after that year. This allows us to measure political outcomes over two election cycles after the layoff and labor market outcomes over up to 4 years after the layoff. [^43] While we must restrict the sample to individuals with unique names for the party membership outcome, we do not impose this restriction when analyzing candidacies, which are less common and thus require larger sample sizes. In addition, we exclude from the sample all workers who were displaced on the first or the last day of each month in their last layoff. Appendix Figure C.1 shows that the number of dismissals is higher at the very beginning and the very end of the month. Dropping these observations ensures that our RD cutoff does not coincide with spikes in the density of dismissals that occur every 30 days, i.e., that are not specific to the 16-month cutoff.
@@
Appendix Figure C.2 shows that, when using the restricted sample, there is no evidence of discontinuity in the density of the running variable at the 16-month cutoff. We also confirm the continuity using the bias-robust test by cattaneo2018; cattaneo2020. In addition, in Appendix Figure C.3 we show that individuals around the cutoff are similar in a rich set of pre-determined covariates, including gender, education, age, earnings, tenure, and sector shares. Overall, these results provide strong support for the continuity assumption required by the RD approach.
@@ 5.2 Results
Results from model (4) are presented in Table 3, and the corresponding RD plots are shown in Figure 8. As a first step, we estimate the effects of UI eligibility on labor market outcomes (Panel A, columns (1) and (2) of Table 3). We find a significant negative effect on employment that amounts to 0.68 fewer months employed over the four years after the layoff. Moreover, labor earnings decrease by 2,192 BRL. The drop in labor earnings is likely compensated by the UI payments that eligible individuals receive. [^44] Britto2022 find that individuals barely meeting the 16-months cutoff receive additional UI benefits of about 2,100 BRL. Thus, the results suggest that UI eligibility leads to a rise in time availability following the job loss without strongly affecting the total available income.
(1) (2) (3) (4) (5) (6) Outcome: Months Labor New party membership (%) Candidacy (%) employed earnings before layoff after layoff before layoff after layoff [A] Full sample $\beta$ -0.679∗∗∗ -2,192.5∗∗∗ -0.002 0.150∗∗ -0.011 0.028 (0.041) (149.9) (0.080) (0.072) (0.013) (0.019) Control mean 26.138 56,281.8 4.330 3.467 0.236 0.449 Relative effect -2.6% -3.9% -0.1% 4.3% -4.6% 6.2% Observations 2,048,674 2,048,674 1,047,683 1,047,683 2,048,674 2,048,674 [B] Low educated (<< 12 years of education) $\beta$ -0.617∗∗∗ -1,822.6∗∗∗ 0.071 0.066 -0.015 -0.006 (0.057) (166.1) (0.121) (0.109) (0.019) (0.026) Control mean 25.029 47,526.6 4.264 3.454 0.221 0.428 Relative effect -2.5% -3.8% 1.7% 1.9% -6.9% -1.4% Observations 986,755 986,755 454,665 454,665 986,755 986,755 [C] High educated ($\geq$ 12 years of education) $\beta$ -0.708∗∗∗ -2,311.2∗∗∗ -0.059 0.216∗∗ -0.007 0.061∗∗ (0.059) (240.6) (0.106) (0.097) (0.019) (0.027) Control mean 27.143 64,213.1 4.381 3.476 0.250 0.468 Relative effect -2.6% -3.6% -1.3% 6.2% -2.7% 13.0% Observations 1,061,919 1,061,919 593,018 593,018 1,061,919 1,061,919 Note: The table presents results from model (4) for the effect of unemployment benefit eligibility on the number of months employed and the total formal labor earnings in the four years after the layoff (columns (1) and (2)), on the likelihood of newly registering as a party member in the two election cycles before and after the layoff (columns (3) and (4)), and on the likelihood of running for local councilor in the two election cycles before and after the layoff (columns (5) and (6)). The sample includes all workers displaced within a bandwidth of 60 days around the cutoff date that determines benefit eligibility, which is 16 months since the last layoff. Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means in columns (3) to (6) have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
A. Full sample Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption B. Low educated (<< 12 years of education) Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption C. High educated ($\geq$ 12 years of education) Months employed New party membership Candidacy Refer to caption Refer to caption Refer to caption Note: The figure plots the average outcomes of workers around the 16-month cutoff date for eligibility for unemployment benefits, for different groups of workers. Outcomes are the number of months employed in the four years after the layoff, the probability of newly registering as a party member in the two election cycles after the layoff (in %), and the probability of running for local councilor in the two election cycles after the layoff (in %). Panel A shows results among all workers in the sample, and Panels B and C distinguish workers with less than vs. at least 12 years of education. Dots show averages for 5-day bins. The lines represent a local linear polynomial smoothing using a 60-day bandwidth, together with 95% confidence intervals.
@@
Given that earlier we have documented differential effects of job loss by education, Panels B and C of Table 3 also show the effects of UI benefits separately for low-educated workers (<< 12 years of schooling) and high-educated workers ($\geq$ 12 years of schooling). We find positive effects on party memberships and candidacies among the highly educated but no significant impact among the low educated. The effects on workers with high levels of education are statistically significant and meaningful in size, indicating a 6.2% increase in the likelihood of becoming a party member and a 13.0% increase in the likelihood of running for office (both relative to the control mean). These findings are consistent with the time-resources mechanism: by reducing reemployment incentives, eligibility for UI benefits frees up time that can be redirected toward political engagement. As suggested by our theoretical framework in Section 2.3, the importance of this channel increases with the expected benefits of political participation. These benefits may vary with education – for example, if more educated individuals have greater chances of electoral success or are more likely to leverage political connections for gaining access to public-sector jobs. Such mechanisms can explain why both job loss and UI eligibility have stronger effects on political entry among the highly educated.
@@ 6.1 Individual-level evidence
To explore the impact of layoffs on political selection, we examine the individual characteristics of candidates and political members who enter politics as a result of being laid off. We use a combination of sample moments in an instrumental variable framework to estimate the average characteristics of three groups: (1) compliers – individuals who enter politics because of the layoff, (2) always-takers – those who enter politics regardless of whether they are laid off or not, and (3) never-takers – those who do not enter politics regardless of their layoff status. To identify compliers, we build on the kappa weighting approach introduced by abadie2003. We estimate the expected value of covariate $X_{i}$ for compliers as
@@
Here, $Y_{i}$ indicates whether an individual enters politics in a post-layoff cycle (either as a party member or a councilor candidate), and $Treat_{i}$ represents their layoff status. We estimate the propensity score, $P(X_{i})=P(Treat_{i}=1|X_{i})$, using the set of all covariates considered in the complier analysis.464646sloczynski2025 provide a comprehensive treatment of various kappa weighting estimators that are implemented in the Stata command kappalate. We use the estimator based on their $\kappa_{i1}$ definition. The authors recommend these kappa weights for applications without always-takers (one-sided noncompliance) to avoid near-zero denominators. Similarly, in our case, the number of always-takers is very small relative to the number of never-takers.
@@
When analyzing the characteristics of political entrants, we focus on measures that capture individual competence, in addition to socio-demographic characteristics. First, we consider years of education as a standard measure of competence. [^47] For Brazil, melo2024 provide evidence that more educated mayors are more effective managers of public resources, as they commit fewer municipal administration infringements. However, they are not found to engage in fewer corrupt practices or improve the provision of local public goods. Second, we examine individuals’ log hourly wage before the layoff, which, following dalbo2013, can be treated as a measure of skill, as the private sector typically rewards highly skilled workers with higher wages. Third, we consider person wage effects from an AKM model that regresses log wages on person fixed effects and firm fixed effects (abowd1999). The person fixed effects measure the wage premia that workers receive independent of the firm that they work at, likely reflecting workers’ ability that is rewarded across all firms. [^48] To our knowledge, we are the first to use AKM effects to study the competence of politicians. The AKM model is fitted on all RAIS contracts in the pre-layoff years 2004 to 2008. Appendix Table D.1 reports a summary of the AKM estimation results. Moreover, we consider residualized AKM person effects, controlling for gender, race, age deciles, education categories, as well as the full set of interactions between these variables. The resulting residuals capture wage premia relative to observably similar workers.494949besley2017 and DalBo2017 also consider residuals from Mincer wage regressions. Our AKM measures additionally purge wages from firm variation in wage-setting, which may not reflect individual ability. To validate our competence measures, we correlate them with local councilors’ misconduct detected during the electoral process and with their electoral performance. Results are reported in Appendix Table D.2 and show that all our competence measures are significantly negatively correlated with the probability of misbehaving in the electoral process and significantly positively related the the probability of being elected. [^50] The misbehavior variable captures a broad range of irregularities detected by the Brazilian Electoral Court during the elections in 2016 and 2020 that resulted in candidacy annulment. These include abuse of power, failure to meet registration requirements, vote buying, prohibited conduct, ineligibility under the “Clean Record Act” (Lei da Ficha Limpa), illicit use of campaign resources, objections to the candidacy, rejection by the party or coalition, and invalidation of the party. Finally, in addition to the competence measures, we also study socio-demographic representation in terms of gender and age.